b. Combine or aggregate the data for the junior and senior years into one crosstabulation as follows: Combined 2-Year Batting A. Fealey Outcome Hit No Hit Total At Bats Calculate each players batting average for the combined two years. E. Janson Junior/Senior Years Allison Fealey Emily Janson Which player should be awarded the scholarship? Select your answerY Explain. The input in the box below will not be graded, but may be reviewed and considered by your instructor c.Are the recommendations you made in parts (a) and (b) consistent? Select your answer- Explain any apparent inconsistencies The input in the box below will not be graded, but may be reviewed and considered by your instructor

those pictures are for the one question, please answer all of them. Thanks!
Answer #1

SOLUTION:

a)

As per above data,

-> Junior :-

Allison Fealey: 15/40 = 0.37

Emily Janson: 70/200 = 0.35

-> Senior :-

Allison Fealey 75/250 = 0.30

Emily Janson 35/120 = 0.29

Allison Fealey should be awarded scholarship since her batting average in both years is higher than Emily Janson.

b)

Combined 2-Year Batting
Outcome A. Fealey E.Janson
Hit 85 110
No Hit 155 260

Total At-Bats 240 370

The batting average of each player for combined two years is:

Combined two Years:

Allison Fealey 85/240 = 0.35

Emily Janson 110/320 = 0.34

Here also Allison Fealey should be awarded scholarship since her batting average in combined years is higher than Emily Janson.

c)

Allison Fealey received scholarship in both cases so the recommendations in parts (a) and (b) is same.